OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - 2y = 6 2x + y = 5 -4x - 2y = 10 -4y = 4 -4x - 5y = -1 -7y = 5 3x + 2y = 6 7x = 12 4x + 2y = 10 8x = 16 @campbell_st @undeadknight26

OpenStudy (campbell_st):

what solution are you given..?

OpenStudy (anonymous):

I dont know thats why im confused

OpenStudy (anonymous):

4x-2y=6 2x+y=5

OpenStudy (campbell_st):

ok... so here is what I would do for this question solve the equations... double the 2nd equation and add the 1st ... this eliminates y 4x - 2y = 6 4x + 2y = 10 -------------- 8x = 16 can you find the value of x..?

OpenStudy (anonymous):

x would be 2?

OpenStudy (campbell_st):

that'c correct... not substitute that into either equation and find y what do you get..?

OpenStudy (anonymous):

8+2y=10 then subract 8 to both sides and get 2y=10 and divide so y equals 5?

OpenStudy (campbell_st):

use the 2nd of the original equations 2x + y = 5 subsitute x = 2 and find y

OpenStudy (campbell_st):

your attempt has an error its 4x - 2y = 6 so 8 - 2y = 6

OpenStudy (anonymous):

4+y=5 y equals 1

OpenStudy (anonymous):

Is that right? @campbell_st

OpenStudy (campbell_st):

that's correct.... so now to find the correct set of equivalent solutions subsitute x = 2 and y = 1 into each pair to find the correct answer